A shopkeeper bought some eggs at 15 cents each. Six of them were broken while the rest were sold at 20 cents each. If she made a profit of $4.80, find how many eggs she bought.
Answer by haileytucki(390)

First, we know that 6 time 15 (the ones broken) = .90 cents.
If she made a profit of 4.80, you have to add the loss of 90cents,
equaling, 5.70 cents.
Now, a little algebra:


.20x =4.80


0.2x=4.8
Divide each term in the equation by 0.2.
(0.2x)/(0.2)=(4.8)/(0.2)
Cancel the common factor of 0.2.
x=(4.8)/(0.2)
Divide 4.8 by 0.2 to get 24.
x=24 are the ones she sold at 20 cent apiece;
you add the 6 she broke= 30 eggs